OpenStudy (anonymous):

On a certain hot summer's day, 660 people used the public swimming pool. The daily prices are $1.25 for children and $2.25 for adults. The receipts for admission totaled $1163.00. How many children and how many adults swam at the public pool that day?

OpenStudy (anonymous):

y + x = 600 1.25x + 2.25y = 1163

OpenStudy (anonymous):

how do I find x and y tho?

OpenStudy (anonymous):

i got;;

OpenStudy (anonymous):

1.25x + 2.25y = 1163 x + y = 600 <--- multiply by 1.25 so you have; 1.25x + 2.25y = 1163 1.25x + 1.25y = 600 -------------------- 1y = 413 adults then plug in 413 into x + y = 600 x + 413 = 600 - 413 - 413 ---------------- x = 187 children

OpenStudy (anonymous):

wow thank you so much!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

and sorry 1 mistake, where u multiply by 1.25 it's 750!

OpenStudy (anonymous):

not 600 anymore.

OpenStudy (anonymous):

337 children********

OpenStudy (anonymous):

omg srry!

OpenStudy (anonymous):

337, cause 750 - 413 adults = 337 kids.